### v3

@@ Appendix D Anatomy of the Coming-Out Event
D.1 Simultaneous Treatments
As described in Section 2.2, the disclosure of Graves’ sexual orientation coincided with the start of LGBT Pride Month. This means that the coming-out event encompasses two treatments occurring at the same time, namely the announcement of Graves’ homosexuality and the introduction of visual and expressive elements in League of Legends that support the LGBT community.454545 See, e.g., [197] for a discussion on “simultaneous treatments” and methodologies for disentangling their effects under a Difference-in-Differences identification strategy.
We recognize the potential influence of LGBT Pride Month on players’ preferences for characters by introducing the binary variable $L_{i}\in\left\{0,1\right\}$ to represent character $i$’s inclusion in the LGB community no later than $T^{pre}+1$. Consequently, we observe three distinct groups of units: the first group includes only Graves, with $C_{i}=L_{i}=1$; the second group includes only Diana, Leona, Nami, and Neeko, with $C_{i}=0$ and $L_{i}=1$; and the third group includes all other characters, with $C_{i}=L_{i}=0$.464646 Neglecting the presence of two simultaneous treatments and treating them as a single treatment does not invalidate the results of Section 4.1. It primarily affects their interpretation, which, without further investigation, could only be attributed to the combined effects of simultaneously receiving both treatments $C_{i}$ and $L_{i}$ - referred to as the coming-out event in the main body of the paper.
To explicitly account for the influence of the two treatments $C_{i}$ and $L_{i}$, we define the potential pick rates as $Y_{i,t}^{c,l}$. Then, for each period $t>T^{pre}$, the effect of the coming-out event on players’ preferences for Graves in (1) corresponds to:
$\tau_{t}=Y_{1,t}^{1,1}-Y_{1,t}^{0,0}$ (D.1)
Equation (D.1) shows why we need to be cautious in interpreting the estimated effects of Section 4.1 as solely stemming from the disclosure of Graves’ sexual orientation. Under an extended version of the SUTVA assumption (see Section D.2), we observe $Y_{1,t}=Y_{1,t}^{1,1}$ for all $t>T^{pre}$, and the estimator in (3) effectively targets the counterfactual series $Y_{1,t}^{0,0}$. Consequently, the estimated effects presented in Section 4.1 encompass the combined impacts of both disclosing Graves’ sexual orientation and his affiliation with the LGB community during LGBT Pride Month. This can be formalized as follows:
@@ D.1 Simultaneous Treatments
D.2 Separating Simultaneous Treatment Effects
@@
Under Assumption 1, we can estimate the counterfactual series $Y_{1,t}^{0,0}$ by constructing a synthetic control unit that approximates the pick rates of Graves before the coming-out event as in Section 3.2. Thus, as shown in (D.2), the challenge in disentangling our causal effects of interest is to estimate $Y_{1,t}^{0,1}$ for $t>T^{pre}$, i.e., how Graves’ pick rates would have evolved if Graves were already part of the LGB community prior to the 2022 LGBT Pride Month.

### v4

@@ D.5 Coming Out versus LGBT Pride Month
F.1 Simultaneous Treatments
As described in Section 2.2, the disclosure of Graves’ sexual orientation coincided with the start of LGBT Pride Month. This means that the coming-out event encompasses two treatments occurring at the same time, namely the announcement of Graves’ homosexuality and the introduction of visual and expressive elements in League of Legends that support the LGBT community.515151 See, e.g., [47] for a discussion on “simultaneous treatments” and methodologies for disentangling their effects under a Difference-in-Differences identification strategy.
We recognize the potential influence of LGBT Pride Month on players’ preferences for characters by introducing the binary variable $L_{i}\in\left\{0,1\right\}$ to represent character $i$’s inclusion in the LGB community no later than $T^{pre}+1$. Consequently, we observe three distinct groups of units: the first group includes only Graves, with $C_{i}=L_{i}=1$; the second group includes only Diana, Leona, Nami, and Neeko, with $C_{i}=0$ and $L_{i}=1$; and the third group includes all other characters, with $C_{i}=L_{i}=0$.525252 Neglecting the presence of two simultaneous treatments and treating them as a single treatment does not invalidate the results of Section 4.1. It primarily affects their interpretation, which, without further investigation, could only be attributed to the combined effects of simultaneously receiving both treatments $C_{i}$ and $L_{i}$ - referred to as the coming-out event in the main body of the paper.
To explicitly account for the influence of the two treatments $C_{i}$ and $L_{i}$, we define the potential pick rates as $Y_{i,t}^{c,l}$. Then, for each period $t>T^{pre}$, the effect of the coming-out event on players’ preferences for Graves in (B.1) corresponds to:
$\tau_{t}=Y_{1,t}^{1,1}-Y_{1,t}^{0,0}$ (F.1)
Equation (F.1) shows why we need to be cautious in interpreting the estimated effects of Section 4.1 as solely stemming from the disclosure of Graves’ sexual orientation. Under an extended version of the SUTVA assumption (see Section F.2), we observe $Y_{1,t}=Y_{1,t}^{1,1}$ for all $t>T^{pre}$, and the estimator in (B.3) effectively targets the counterfactual series $Y_{1,t}^{0,0}$. Consequently, the estimated effects presented in Section 4.1 encompass the combined impacts of both disclosing Graves’ sexual orientation and his affiliation with the LGB community during LGBT Pride Month. This can be formalized as follows:
@@ F.1 Simultaneous Treatments
F.2 Separating Simultaneous Treatment Effects
@@
Under Assumption 1, we can estimate the counterfactual series $Y_{1,t}^{0,0}$ by constructing a synthetic control unit that approximates the pick rates of Graves before the coming-out event as in Section B. Thus, as shown in (F.2), the challenge in disentangling our causal effects of interest is to estimate $Y_{1,t}^{0,1}$ for $t>T^{pre}$, i.e., how Graves’ pick rates would have evolved if Graves were already part of the LGB community prior to the 2022 LGBT Pride Month.